# Two-particle calculations with quantics tensor trains - solving the parquet equations

### v1

@@ Abstract
We present the first application of quantics tensor trains (QTTs) and tensor cross interpolation (TCI) to the solution of a full set of self-consistent equations for multivariate functions, the so-called parquet equations. We show that the steps needed to evaluate the equations (Bethe–Salpeter equations, parquet equation and Schwinger–Dyson equation) can be decomposed into basic operations on the QTT-TCI (QTCI) compressed objects. The repeated application of these operations does not lead to the loss of accuracy beyond a specified tolerance and the iterative scheme converges even for numerically demanding parameters. As examples we take the Hubbard model in the atomic limit and the single impurity Anderson model, where the basic objects in parquet equations, the two-particle vertices, depend on three frequencies, but not on momenta. The results show that this approach is able to overcome major computational bottlenecks of standard numerical methods. The applied methods allow for an exponential increase of the number of grid points included in the calculations leading to an exponentially improving computational error for a linear increase in computational cost.
@@ I Introduction
The quantics tensor train (QTT) representation of multivariate functions has already been around for a decade or so [17, 18, 19, 20], but it was only recently applied to various fields of natural science such as turbulence [21, 22, 23], plasma physics [24], quantum chemistry [25], and quantum field theory of the many electron problem [26]. For quantum field theories, the QTT representation provides a compact representation of the space-time dependence of the correlation functions [26]. First many-body calculations of Feynman diagrams with the QTT representation in imaginary time [27] and in nonequilibrium [28] already show the potential of the method. A very favorable scaling of the QTT representation with temperature is conjectured [29]. In parallel, the tensor cross interpolation (TCI) method was applied to evaluations of diagrams in many-body physics [30, 31, 32, 33]. TCI can be combined with the quantics tensor train representation to form QTT+TCI=QTCI, a powerful approach with diverse applications [34].
@@
In order to apply QTCI to parquet equations, we break down these equations (Bethe–Salpeter equation, parquet equation and Schwinger–Dyson equation) into basic operations on QTTs, represented by matrix product operators (MPOs). We use MPO-MPO contractions for matrix and elementwise multiplications and construct a new MPO for affine transformations that are needed to perform channel transformations (variable shifts) occurring in the parquet equation. Our approach scales as $\mathcal{O}({D_{\mathrm{max}}}^{4}R)$, with maximum bond dimension of ${D_{\mathrm{max}}}$ and grid size $2^{3R}$. The computational cost is only logarithmic in grid size and the main bottleneck is shifted to the maximum bond dimension. We have verified this scaling in two benchmark models: the Hubbard atom and the single impurity Anderson model (SIAM). In both models, the two-particle vertices are fully dynamical (dependent on three frequencies) but local (independent of momentum). In both cases, we empirically find that an overall accuracy of $10^{-4}$ of the full self-consistent solution can be achieved with a bond dimension up to 200 even for challenging parameters close to a divergence line in the Hubbard atom.
@@
This paper is organized as follows. In Sec. II we introduce the concrete Hamiltonians (Hubbard atom and single impurity Anderson model) for which we will present the results. Further, in Sec. III we first recall definitions of one- and two-particle Green’s functions and vertices and set the notation used in the paper. Then we provide in detail the full set of parquet equations that we solve. Additional information on the equations and notation is also provided in App. A. In IV we introduce quantics tensor trains, the tensor cross interpolation method and matrix product operators. These techniques are used to construct efficient implementations of the parquet equations in Sec. V and Apps. B-C. We also provide in this section (Sec. V) results for the compression of the vertices and scaling of the bond dimension for each of the operations needed to complete one loop of parquet equations. More details and additional plots can be found in Appendices D-E. Next, in Sec. VI, we show results for the full self-consistent iterative scheme and its technical limitations (with details also in App. F). In the last section VII, we conclude and provide outlook.
@@ II Models
In the current work we focus on the solution of equations for two-particle vertices in the (Matsubara) frequency space. Although in general the vertices are also dependent on momentum and orbital degrees of freedom, we limit ourselves to the simplest case and consider only the full frequency dependence of the vertices. We present results for two benchmark models: the Hubbard atom, where exact analytical expressions for the vertex functions are known [35] and the single impurity Anderson model [36], where high-quality numerical data is available. The treatment of the frequency dependence of vertices presented in this work can be directly extended to models with additional orbital and momentum dependencies.
@@ II.2 Single-impurity Anderson model
where the impurity is described by the fermionic annihilation (creation) operators $\hat{d}_{\sigma}^{(\dagger)}$, the number operator $\hat{n}_{d,\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$, the impurity one-particle energy level $\varepsilon_{d}$ and the onsite repulsion $U$. The bath is described by the kinetic term $\hat{c}_{\bm{k},\sigma}^{(\dagger)}$ with one-particle energies $\varepsilon_{\bm{k}}$. The hybridization between the impurity and the bath is given by $V_{\bm{k}}$. The bath parameters can be put together into a frequency dependent hybridization function:
@@ III.1 One-particle quantities
where $\Delta(\nu)$ vanishes for the Hubbard atom, $\Delta(\nu)$ equals Eq. (4) for the single-impurity Anderson model, and in both models we have set the chemical potential and the one-particle energy level $\varepsilon_{d}$ of the impurity to half filling.
@@ III.2 Two-particle quantities
In the above definition of the Fourier transform, the two-particle Green’s function is dependent on three fermionic Matsubara frequencies $\nu_{1},\nu_{2},\nu_{3}$. In the context of Bethe–Salpeter equations (defined below in Sec. III.3), it is more convenient to parameterize two-particle quantities as a function of two fermionic frequencies $\nu,\nu^{\prime}$ and one bosonic Matsubara frequency $\omega=\frac{2n\pi}{\beta},n\in\mathbb{Z}$. There are three important conventions of this parametrization: the particle-hole (ph) channel notation, where $\omega=\nu_{1}+\nu_{2}$; the particle-particle (pp) channel notation, where $\omega=\nu_{1}+\nu_{3}$, and the transversal particle-hole ($\overline{ph}$) channel notation, where $\omega=\nu_{2}+\nu_{3}$. In the parquet approach, it is necessary to transform between these conventions using so-called channel transformations outlined in App. A. The reason, as we will see in Sec. III.4, is the parquet equation which mixes vertex functions that are represented in different frequency channel parametrizations.
@@
Apart from channels stemming from different frequency parametrizations (ph, pp, and $\overline{ph}$), it is convenient to introduce also linear combinations of spin components. The following spin combinations will be used for vertices in the ph frequency channel:
@@
The same vertex $F$ can be represented in the particle-particle channel frequency parametrization: $F^{pp}$ (see Appendix A for details). In the pp channel, the convenient spin combinations are the following:
$F_{s}$ $=F^{pp}_{\uparrow\uparrow}-F^{pp}_{\uparrow\downarrow},$
$F_{t}$ $=F^{pp}_{\uparrow\uparrow},$ (11)
physically corresponding to the singlet (s) and triplet (t) spin components. In the following we will predominantly use the spin component notation, i.e. $d/m/s/t$, assuming that the $d$ or $m$ spin components are always in the ph frequency channel notation and the $s$ or $t$ spin components are always in the pp frequency channel notation.
@@ III.3 Bethe–Salpeter equations
$F_{d}^{\nu\nu^{\prime}\omega}=\Gamma_{d}^{\nu\nu^{\prime}\omega}-\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}\Gamma_{d}^{\nu\nu_{1}\omega}\chi_{0,ph}^{\nu_{1}\nu_{2}\omega}F_{d}^{\nu_{2}\nu^{\prime}\omega},$ (12a)
$F_{m}^{\nu\nu^{\prime}\omega}=\Gamma_{m}^{\nu\nu^{\prime}\omega}-\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}\Gamma_{m}^{\nu\nu_{1}\omega}\chi_{0,ph}^{\nu_{1}\nu_{2}\omega}F_{m}^{\nu_{2}\nu^{\prime}\omega};$ (12b)
@@
$F_{s}^{\nu\nu^{\prime}\omega}=\Gamma_{s}^{\nu\nu^{\prime}\omega}\pm\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}F_{s}^{\nu\nu_{1}\omega}\chi_{0,pp}^{\nu_{1}\nu_{2}\omega}\Gamma_{s}^{\nu_{2}\nu^{\prime}\omega},$ (12c)
$F_{t}^{\nu\nu^{\prime}\omega}=\Gamma_{t}^{\nu\nu^{\prime}\omega}\pm\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}F_{t}^{\nu\nu_{1}\omega}\chi_{0,pp}^{\nu_{1}\nu_{2}\omega}\Gamma_{t}^{\nu_{2}\nu^{\prime}\omega}.$ (12d)
@@
$\chi_{0,ph}^{\nu\nu^{\prime}\omega}=-\beta G(\nu)G(\nu+\omega)\delta_{\nu\nu^{\prime}},$ (14a)
$\chi_{0,pp}^{\nu\nu^{\prime}\omega}=-\frac{\beta}{2}G(\nu)G(-\nu-\omega)\delta_{\nu\nu^{\prime}}.$ (14b)
@@ III.4 Parquet equation
Through Eqs. (12)-(13) we defined reducible vertices $\Phi_{d/m}$ and $\Phi_{s/t}$ in ph and pp channels, respectively. These vertices correspond to different physical processes that happen in the ph and pp scattering channels and are generated by the BSEs (12). The parquet equation mixes these processes, allowing for balance between contributions generated by the all of the BSEs (12). In a simplified way, the parquet equation can be represented as the following sum of terms:
$F=\Lambda+\Phi_{ph}+\Phi_{\overline{ph}}+\Phi_{pp},$ (15)
where $\Phi_{ph}$ denotes contributions coming from $\Phi_{d}$ or $\Phi_{m}$, $\Phi_{\overline{ph}}$ contributions coming from $\Phi_{d/m}$, but in the ${\overline{ph}}$ frequency parametrization, and $\Phi_{pp}$ contributions from $\Phi_{s}$ or $\Phi_{t}$ (more details can be found in App. A or in Ref. 6). The last summand, $\Lambda$, contains so-called fully two-particle irreducible diagrams, i.e., contributions which cannot be generated by the two-particle BSEs.
@@
$F_{d}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{d}^{\nu\nu^{\prime}\omega}+\Phi_{d}^{\nu\nu^{\prime}\omega}-\frac{1}{2}\Phi_{d}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}-\frac{3}{2}\Phi_{m}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}$
$\quad+\frac{1}{2}\Phi_{s}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}+\frac{3}{2}\Phi_{t}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})},$ (16)
where $\Lambda_{d}$ is the density component of the fully irreducible vertex. This vertex cannot be obtained from the BSEs and has to be provided from outside the parquet scheme. In the examples presented in Sec. VI, we either use the exact expression (it is known for the Hubbard atom) or we use a weak coupling approximation for it. Eq. (15) can be used to generate equations analogous to Eq. (16) for $F_{m}$, $F_{s}$, and $F_{t}$. We provide them explicitly in App. A.
@@ III.6 Iterative parquet scheme
Figure 1: Iterative parquet scheme
@@ IV Quantics Tensor trains
Numerical solution of the iterative parquet scheme introduced in the previous section suffers from the curse of dimensionality: the vertex functions have multiple frequency and in general also momentum arguments. Discretizing these multivariate functions on naive grids requires a number of grid points that grows exponentially with the number of function arguments, which therefore becomes very expensive already for a moderate number of grid points in each argument. The solution to this problem that we propose is (i) to separate each variable dependence into a set of binary numbers (hence ‘quantics’) corresponding to different length/energy scales; (ii) to factorize the dependence on each argument at each length/energy scale into a tensor train (TT), also known as matrix product state (MPS). If the problem has some kind of scale separation, the resulting quantics tensor train (QTT) is expected to have a small maximum bond dimension. Since this is the case in many physical problems, such an approach is potentially very powerful. It has already been shown to reduce computational costs significantly in several applications with high-dimensional functions [26, 34, 17, 38, 18].
In this section we introduce the definition of the QTT representation, then present a method for efficient compression of multivariate functions into a QTT, namely the tensor cross interpolation (TCI). Finally, we also introduce matrix product operators (MPOs) that are needed for computations with the QTTs. These methods are valid for any multivariate function and not specific to two-particle vertices.

### v2

@@ Abstract
We present the first application of quantics tensor trains (QTTs) and tensor cross interpolation (TCI) to the solution of a full set of self-consistent equations for multivariate functions, the so-called parquet equations. We show that the steps needed to evaluate the equations (Bethe–Salpeter equations, parquet equation and Schwinger–Dyson equation) can be decomposed into basic operations on the QTT-TCI (QTCI) compressed objects. The repeated application of these operations does not lead to a loss of accuracy beyond a specified tolerance and the iterative scheme converges even for numerically demanding parameters. As examples we take the Hubbard model in the atomic limit and the single impurity Anderson model, where the basic objects in parquet equations, the two-particle vertices, depend on three frequencies, but not on momenta. The results show that this approach is able to overcome major computational bottlenecks of standard numerical methods. The applied methods allow for an exponential increase of the number of grid points included in the calculations leading to an exponentially improving computational error for a linear increase in computational cost.
@@ I Introduction
The quantics tensor train (QTT) representation of multivariate functions has already been around for a decade or so [17, 18, 19, 20], but it was only recently applied to various fields of natural science such as turbulence [21, 22, 23, 24, 25], plasma physics [26], quantum chemistry [27], and quantum field theory of the many electron problem [28]. For quantum field theories, the QTT representation provides a compact representation of the space-time dependence of the correlation functions [28]. First many-body calculations of Feynman diagrams with the QTT representation in imaginary time [29] and in nonequilibrium [30] already show the potential of the method. A very favorable scaling of the QTT representation with temperature has been conjectured in Ref. 31. In parallel, the tensor cross interpolation (TCI) method was applied to evaluations of diagrams in many-body physics [32, 33, 34, 35]. TCI can be combined with the quantics tensor train representation to form QTT+TCI=QTCI, a powerful approach with diverse applications [36].
@@
In order to apply QTCI to parquet equations, we break down these equations (Bethe–Salpeter equation, parquet equation and Schwinger–Dyson equation) into basic operations on QTTs, represented by matrix product operators (MPOs). We use MPO-MPO contractions for matrix and elementwise multiplications and construct a new MPO for affine transformations that are needed to perform channel transformations (variable shifts) occurring in the parquet equation. Our approach scales as $\mathcal{O}(D_{\mathrm{max}}^{4}R)$, with maximum bond dimension of $D_{\mathrm{max}}$ and grid size $2^{3R}$. The computational cost is only logarithmic in grid size and the main bottleneck is shifted to the maximum bond dimension. We have verified this scaling in two benchmark models: the Hubbard atom and the single impurity Anderson model (SIAM). In both models, the two-particle vertices are fully dynamical (dependent on three frequencies) but local (independent of momentum). In both cases, we empirically find that an overall accuracy $<10^{-3}$ of the full self-consistent solution can be achieved with a bond dimension up to 200 even for challenging parameters close to a divergence line in the Hubbard atom.
@@
This paper is organized as follows. In Sec. II we introduce the concrete Hamiltonians (Hubbard atom and single impurity Anderson model) for which we will present the results. Further, in Sec. III we first recall definitions of one- and two-particle Green’s functions and vertices and set the notation used in the paper. Then we provide in detail the full set of parquet equations that we solve. Additional information on the equations and notation is also provided in App. A. In IV we introduce quantics tensor trains, the tensor cross interpolation method and matrix product operators. These techniques are used to construct efficient implementations of the parquet equations in Sec. V and Apps. B-C. We also provide results for the compression of the vertices and scaling of the bond dimension for each of the operations needed to complete one loop of parquet equations in Sec. V. More details and additional plots can be found in Appendices D-E. Next, in Sec. VI, we show results for the full self-consistent iterative scheme and its technical limitations (with details also in App. F). In the last section VII, we conclude and provide outlook.
@@ II Models
In the current work we focus on the solution of equations for two-particle vertices in the (Matsubara) frequency space. Although in general the vertices are also dependent on momentum and orbital degrees of freedom, we limit ourselves to simple models for which the vertices depend only on frequency but not on momentum. We present results for two benchmark models: the Hubbard atom, where exact analytical expressions for the vertex functions are known [37] and the single impurity Anderson model [38], where high-quality numerical data is available  [3]. The treatment of the frequency dependence of vertices presented in this work can be directly extended to models with additional orbital and momentum dependencies.
@@ II.2 Single-impurity Anderson model
where the impurity is described by the fermionic annihilation (creation) operators $\hat{d}_{\sigma}^{(\dagger)}$, the number operator $\hat{n}_{d,\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$, the impurity one-particle energy level $\varepsilon_{d}$ and the onsite repulsion $U$. The bath is described by the kinetic term $\hat{c}_{\bm{k},\sigma}^{(\dagger)}$ with one-particle energies $\varepsilon_{\bm{k}}$. The hybridization between the impurity and the bath is given by $V_{\bm{k}}$. The bath parameters jointly determine the frequency dependent hybridization function:
@@ III.1 One-particle quantities
where we have set the following model-dependent parameters to values corresponding to half filling: For the Hubbard atom, $\Delta(\nu)$ vanishes, and the chemical potential is $\mu=U/2$. For the single-impurity Anderson model, we set the chemical potential to $\mu=0$ and the one-particle energy level to $\epsilon_{d}=-U/2$.
@@ III.2 Two-particle quantities
In the above definition of the Fourier transform, the two-particle Green’s function is dependent on three fermionic Matsubara frequencies $\nu_{1},\nu_{2},\nu_{3}$. In the context of Bethe–Salpeter equations (defined below in Sec. III.3), it is more convenient to parameterize two-particle quantities as a function of two fermionic frequencies $\nu,\nu^{\prime}$ and one bosonic Matsubara frequency $\omega=\frac{2n\pi}{\beta},n\in\mathbb{Z}$. There are three important conventions of this parametrization: the particle-hole (ph) channel notation, where $\omega=\nu_{1}+\nu_{2}$; the particle-particle (pp) channel notation, where $\omega=\nu_{1}+\nu_{3}$, and the transversal particle-hole ($\overline{\textit{ph}}$) channel notation, where $\omega=\nu_{2}+\nu_{3}$. In the parquet approach, it is necessary to transform between these conventions using so-called channel transformations outlined in App. A. The reason, as we will see in Sec. III.4, is that the parquet equation mixes vertex functions that are represented in different frequency channel parametrizations.
@@
Apart from channels stemming from different frequency parametrizations (ph, pp, and $\overline{\textit{ph}}$), it is convenient to introduce also linear combinations of spin components. The following spin combinations will be used for vertices in the ph frequency channel:
@@
The same vertex $F$ can be represented in the particle-particle channel frequency parametrization: $F^{\textit{pp}}$ (see Appendix A for details). In the pp channel, the convenient spin combinations are the following:
$F_{s}$ $=F^{\textit{pp}}_{\uparrow\uparrow}-F^{\textit{pp}}_{\uparrow\downarrow},$
$F_{t}$ $=F^{\textit{pp}}_{\uparrow\uparrow},$ (11)
physically corresponding to the singlet (s) and triplet (t) spin components. In the following, we will predominantly use the spin component notation, i.e. $d/m/s/t$, assuming that the $d$ or $m$ spin components are always in the ph frequency channel notation and the $s$ or $t$ spin components are always in the pp frequency channel notation.
@@ III.3 Bethe–Salpeter equations
$F_{d}^{\nu\nu^{\prime}\omega}=\Gamma_{d}^{\nu\nu^{\prime}\omega}-\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}\Gamma_{d}^{\nu\nu_{1}\omega}\chi_{0,\textit{ph}}^{\nu_{1}\nu_{2}\omega}F_{d}^{\nu_{2}\nu^{\prime}\omega},$ (12a)
$F_{m}^{\nu\nu^{\prime}\omega}=\Gamma_{m}^{\nu\nu^{\prime}\omega}-\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}\Gamma_{m}^{\nu\nu_{1}\omega}\chi_{0,\textit{ph}}^{\nu_{1}\nu_{2}\omega}F_{m}^{\nu_{2}\nu^{\prime}\omega};$ (12b)
@@
$F_{s}^{\nu\nu^{\prime}\omega}=\Gamma_{s}^{\nu\nu^{\prime}\omega}+\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}F_{s}^{\nu\nu_{1}\omega}\chi_{0,\textit{pp}}^{\nu_{1}\nu_{2}\omega}\Gamma_{s}^{\nu_{2}\nu^{\prime}\omega},$ (12c)
$F_{t}^{\nu\nu^{\prime}\omega}=\Gamma_{t}^{\nu\nu^{\prime}\omega}-\frac{1}{\beta^{2}}\sum_{\nu_{1}\nu_{2}}F_{t}^{\nu\nu_{1}\omega}\chi_{0,\textit{pp}}^{\nu_{1}\nu_{2}\omega}\Gamma_{t}^{\nu_{2}\nu^{\prime}\omega}.$ (12d)
@@
$\chi_{0,\textit{ph}}^{\nu\nu^{\prime}\omega}=-\beta G(\nu)G(\nu+\omega)\delta_{\nu\nu^{\prime}},$ (14a)
$\chi_{0,\textit{pp}}^{\nu\nu^{\prime}\omega}=-\frac{\beta}{2}G(\nu)G(-\nu-\omega)\delta_{\nu\nu^{\prime}}.$ (14b)
@@ III.4 Parquet equation
Through Eqs. (12)-(13) we defined reducible vertices $\Phi_{d/m}$ and $\Phi_{s/t}$ in ph and pp  channels, respectively. These vertices correspond to different physical processes that happen in the ph and pp scattering channels and are generated by the BSEs (12). The parquet equation mixes these processes, allowing for balance between contributions generated by all of the BSEs (12). In a simplified way, the parquet equation can be represented as the following sum of terms:
$F=\Lambda+\Phi_{\textit{ph}}+\Phi_{\overline{\textit{ph}}}+\Phi_{\textit{pp}},$ (15)
where $\Phi_{\textit{ph}}$ denotes contributions coming from $\Phi_{d}$ or $\Phi_{m}$, $\Phi_{\overline{\textit{ph}}}$ contributions coming from $\Phi_{d/m}$, but in the $\overline{\textit{ph}}$ frequency parametrization, and $\Phi_{\textit{pp}}$ contributions from $\Phi_{s}$ or $\Phi_{t}$ (more details can be found in App. A or in Ref. 6). The first summand, $\Lambda$, contains so-called fully two-particle irreducible diagrams, i.e., contributions which cannot be generated by the two-particle BSEs.
@@
$F_{d}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{d}^{\nu\nu^{\prime}\omega}+\Phi_{d}^{\nu\nu^{\prime}\omega}-\tfrac{1}{2}\Phi_{d}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}$
$\quad-\tfrac{3}{2}\Phi_{m}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}+\tfrac{1}{2}\Phi_{s}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}$
$\quad+\tfrac{3}{2}\Phi_{t}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}\,,$ (16)
where $\Lambda_{d}$ is the density component of the fully irreducible vertex. $\Lambda_{d}$ cannot be obtained from the BSEs and has to be provided from outside the parquet scheme. In the examples presented in Sec. VI, we either use the exact expression (it is known for the Hubbard atom) or we use a weak coupling approximation for it. Eq. (15) can be used to generate equations analogous to Eq. (16) for $F_{m}$, $F_{s}$, and $F_{t}$. We provide them explicitly in App. A.
@@ III.6 Iterative parquet scheme
Figure 1: Iterative parquet scheme.
@@ IV Quantics Tensor trains
Numerical solution of the iterative parquet scheme introduced in the previous section suffers from the curse of dimensionality: the vertex functions have multiple frequency (and in general also momentum) arguments. Discretizing these multivariate functions on naive grids requires a number of grid points that grows exponentially with the number of function arguments, which therefore becomes very expensive already for a moderate number of grid points in each argument. The solution to this problem that we propose is (i) to represent each variable through a set of binary numbers (hence ‘quantics’) corresponding to different length/energy scales; (ii) to factorize the dependence on each argument at each length/energy scale into a tensor train (TT), also known as matrix product state (MPS). If the problem has some kind of scale separation, the resulting quantics tensor train (QTT) is expected to have a small maximum bond dimension. Since this is the case in many physical problems, such an approach is potentially very powerful. It has already been shown to reduce computational costs significantly in several applications with high-dimensional functions [28, 36, 17, 40, 18].
In this section, we introduce the definition of the QTT representation, then present a method for efficient compression of multivariate functions into a QTT, namely the tensor cross interpolation (TCI). Finally, we also introduce matrix product operators (MPOs) that are needed for computations with the QTTs. These methods are valid for any multivariate function and not specific to two-particle vertices.